OpenStudy (anonymous):

Dot product, determine whether the three points are vertices of a triangle A=(-2,3), B=(3,1), C=(-4,-1)

OpenStudy (anonymous):

Dot product is also known as scalar product

OpenStudy (anonymous):



OpenStudy (anonymous):

How did you find the vertices with the dot product?

OpenStudy (anonymous):

the question gives you the verticies. So i made a small plot to show approximately where they would be.

OpenStudy (anonymous):

What is the dot product equation?

OpenStudy (anonymous):

for two points it is A = (x1, y1) B = (x2, y2) x1 * x2 + y1 * y2 it is safe to assume that with all three points, you would multiply all three x values. so it would look something like this A=(-2,3), B=(3,1), C=(-4,-1) Dot Product = -2*3*-4 + 3*1*-1

OpenStudy (anonymous):

24 - 3 = 21

OpenStudy (anonymous):

where did you get 24-3 from

OpenStudy (anonymous):

im not too sure where to go with this. its combining the dot product that we found.

OpenStudy (anonymous):

I was thinking maybe we need to put the current vertices into vectors since we are looking for the dot product.

OpenStudy (anonymous):

lets do our dot products with only two points at a time A=(-2,3), B=(3,1), C=(-4,-1) A dot B = -2 * 3 + 3 * 1 = -3 A dot C = -4*-2 + 3*-1 = 5 B dot C = -4 * 3 + 1 * -1 = -13 AB = -3, AC = 5 BC = -13

OpenStudy (anonymous):

I got that too, but when I worked out into a^2+b^2=c^2 it doesn't equal a right triangle

OpenStudy (anonymous):

not everything has to be a right triangle. this is not a triangle. The longest side is longer than the sum of the two smaller sides, so the triangle would not occur.